OpenStudy (anonymous):

y varies indirectly as twice x. When x = 1, y = 3. Find y when x = 4.

OpenStudy (anonymous):

12

OpenStudy (anonymous):

it has to be a fraction

OpenStudy (anonymous):

oh then idk

OpenStudy (anonymous):

lol its cool

OpenStudy (anonymous):

how bout 1/3=4/x and then solve

OpenStudy (anonymous):

its 3/4

OpenStudy (anonymous):

OH 3/4=x/1 makes sense